OpenStudy (anonymous):

∫ (x^2)/(2+x^3) dx I have in my own work: ∫ x^2(2+x^3)^-1 dx with u=2+x^3 du=3x^2 dx=du/3x^2 So ∫ x^2 (u)^-1 du/(3x^2) ∫ 1/3 (u)^-1 And that's where I get lost b/c the anti derivative of (u)^-1 is (u)^0 which can't be right. Please show your steps.

OpenStudy (lalaly):

your right but the last step\[\int\limits{\frac{1}{3u}du}=\frac{1}{3} \int\limits{\frac{du}{u}}=\frac{1}{3}\ln(u)\]

OpenStudy (lalaly):

because integral of 1/x =ln(x) not x^0

hartnn (hartnn):

lastly don't forget to re-substitute u= 2+x^3 in 1/3 ln u+c

OpenStudy (anonymous):

So my anti derivative of (u)^-1 would just be u/1? or ln(u)?

hartnn (hartnn):

antiderivative of u^-1 is ln u+c not u/1 or u^0